OpenStudy (puppysized):

A skydiver jumps from a plane that is 5,000 feet above the ground. He opens his parachute when he is 1,500 feet above the ground. a) Write an equation that gives the height in feet of the skydiver above the ground in seconds. b) For how many seconds does the skydiver fall before opening his parachute? Round your answer to the nearest hundredth. For a bit of background: I'm tutoring someone in math, and this is a problem we have to finish. It's in Algebra II, a section on quadratics and factoring, and this is all we were given to solve it. It probably requires some physics to solve...help?

OpenStudy (australopithecus):

a) Since it is algebra 2 then I assume you need no calculus You would use kinematics equations for this problem this one in particular, its derivation requires integration and since it is algebra 2 I dont think you need to provide it. Also, since nothing is given and I dont expect wind resistance to be factored in or drag or whatever (also because they make the calculations much more complicated) you just need this one formula The change in x equals x2 - x1 = (1/2)a*t^2 x2 = final position x1 = initial position Just used the acceleration of gravity g = a = 32.19ft/s^2 Rearrange the equation to get t b) Just use the formula (rearranged to equal t) x2 - x1 = (1/2)a*t^2 this part gives you x2 and x1 so you have no unknowns except for t